# Xu Duan, Lei Sun [^2] Corresponding author. , and Xiangxiang Chu Amap, Alibaba Group

### v2

@@ 3.4 Unified Control Alignment
Benefiting from the architectural simplicity of SCALAR, extending it to manage multiple conditions simultaneously requires addressing a key challenge: the control features extracted for different modalities (e.g., Canny edges, depth maps) reside in distinct and potentially incompatible feature spaces. As shown in Figure 4, we propose a unified version, SCALAR-Uni. The core idea is to map disparate control features into a common, modality-agnostic latent space. We utilize the image feature space itself as the target for this alignment, as it offers a rich and universal representation of visual concepts. We enforce this alignment by introducing an auxiliary loss term during training, which minimizes the L2 distance between the projected control representations and the corresponding image features:
@@
Figure 5: Visual results generated by SCALAR.
Figure 6: Visualization of controllable image synthesis with SCALAR by changing class labels (e.g., “polecat” → “goldfish”).
Figure 7: Visualization of zero-shot controllable image synthesis with SCALAR, including inpainting and outpainting. From top to bottom, Depth, Sketch, Normal, Canny, and HED are used as control image inputs.
@@ 4.1 Experimental Setup
Dataset. We conduct experiments on the ImageNet-256 [5], using all 50K images from the validation set for evaluation. To assess the controllable generation capability of our model, we consider five types of conditions: Canny [4], Depth [35], Normal [50], HED [52], and Sketch [42].
Training Details. We extract four features from the Image Encoder $\mathcal{E}$ at the layers specified by the index set =$\mathcal{I}=\{d-1-k\cdot\lfloor d/4\rfloor\mid k\in\{0,1,2,3\}\}$. We use DINOv2 [29] as the Image Encoder for its strong representations and scalability. As shown in Table 1, we use pretrained VARs with depths of 12, 16, 20, and 24, adopting DINOv2-S for 12 and DINOv2-B otherwise. Following previous works [58, 18], the control block is zero-initialized. Our SCALAR and SCALAR-Uni are trained for 10 epochs with the AdamW optimizer on 8 NVIDIA H20 GPUs. For SCALAR-Uni, we sample different control conditions with equal probability.
Evaluation Metrics. We mainly employ two metrics: conditional consistency and image generation quality. We evaluate the conditional consistency by calculating the similarity between the input condition images and the extracted condition images from the generated images. Specifically, we use F1-Score to assess the similarity of Canny and Sketch, Root Mean Square Error (RMSE) for Normal and Depth, and Structural Similarity Index Measure (SSIM) for HED. For image generation quality, we use Fréchet Inception Distance (FID) [10] and Inception Score (IS) [38].
4.2 Experimental Results
We evaluate the controllable generation performance of our proposed SCALAR and SCALAR-Uni on the ImageNet [5]. As shown in Table 1, we compute both conditional consistency and image generation quality for the images generated by SCALAR and SCALAR-Uni. Furthermore, we compare our approach with existing controllable generation methods, including diffusion-based models [58, 26], raster-scan AR models [21], and VAR-based methods [20, 53].
@@ 4.2 Experimental Results
Figure 8: Visual results generated by our unified multi-condition control method SCALAR-Uni under varying control conditions.
Results for SCALAR-Uni. Compared to other SoTA methods [58, 26, 21, 20, 53], SCALAR-Uni demonstrates clear advantages in image generation quality while maintaining high conditional consistency. When compared to SCALAR, SCALAR-Uni shows a slight drop in both generation quality and consistency, which may be attributed to the need to accommodate multiple control conditions, thereby reducing the effective training data per condition. Overall, SCALAR-Uni exhibits strong performance and generalization as a unified framework for multi-condition controllable generation. More visualization details can be found in Figure 8.
Figure 9: Visualization of zero-shot hybrid controllable image synthesis with SCALAR-Uni, combining two control types (e.g., Depth+Sketch) to generate images with both characteristics. Different class-condition injections are applied for each control pair.
@@ 4.3 Ablation Study
Ablations on Image Encoder $\mathcal{E}$. In Table 2, we conduct experiments using different image encoders (or pretraining schemes) towards different controls on ImageNet [5], including Canny and Depth. Firstly, unlike previous approaches [21, 53], SCALAR and SCALAR-Uni adopt frozen image encoders [6, 29, 14] to preserve the robust features obtained from large-scale self-supervised pretraining. In addition, inspired by prior work [54, 3, 55], we extract multi-layer features including intermediate layers, instead of relying solely on the final output, which improves image generation quality while maintaining control consistency. As shown in Table 2 (c) to (g), DINOv2 [29] outperforms other pretrained vision models such as ViT [6] and SAM [14]. Furthermore, comparing (d), (e), (c), and (g) indicates that increasing the scale of the image encoder significantly boosts performance under the same backbone. Considering the trade-off between parameter size and effectiveness, we use DINOv2-s for the d12 architecture and DINOv2-b for all other settings. Finally, the comparison from (g) to (j) demonstrates that deeper VAR networks lead to further improvements in both generation quality and control accuracy, suggesting that deeper models can better exploit visual representations.
@@
Ablations on Unified Control Alignment. We evaluate our SCALAR-Uni under different depths (d12, d16, d20, and d24) on ImageNet [5]. While the model is trained with five types of control conditions, the effectiveness of Unified Control Alignment is validated only on Canny and Depth. As shown in Table 3, without the Unified Control Alignment, SCALAR-Uni is equivalent to SCALAR trained with mixed supervision across the five control types. Thanks to the strong general visual representation capability of DINOv2, SCALAR can effectively capture the differences across modalities and achieve preliminary alignment of multi-modal control signals. After incorporating Unified Control Alignment, SCALAR-Uni consistently outperforms SCALAR across different network depths in both image generation quality and control consistency. Results in Table 3 demonstrate that SCALAR-Uni, constructed by integrating SCALAR with Unified Control Alignment, offers a simple and effective solution for controllable image generation under diverse conditions.
@@ 5 Conclusion and Future Work
Future Work. (1) To align with the next-scale paradigm in VAR, we directly inject control features extracted from the image encoder into corresponding scales after resizing them via bilinear interpolation. While simple to implement, this approach is relatively coarse and lacks fine-grained semantic alignment, which may limit the further improvement of SCALAR’s controllability. This issue is worth further study. (2) Due to resource constraints, our SCALAR is currently evaluated only on c2i tasks using the ImageNet, where it demonstrates strong performance in both generation quality and controllability. In future work, we plan to extend SCALAR to controllable t2i generation tasks and validate its effectiveness on diverse text-image paired datasets. (3) SCALAR successfully combines the powerful generation capabilities of VAR with a simple and effective control injection mechanism. We are excited to inspire related tasks (e.g., image editing).
Table 2: Ablation of Image Encoder $\mathcal{E}$ and the backbone depth of VAR. xx cells mark Image Encoder parameters comparable to CAR [53] / ControlAR [21], with a smaller backbone.

### v3

@@ 3.4 Unified Control Alignment
Benefiting from the architectural simplicity of SCALAR, extending it to manage multiple conditions simultaneously only requires addressing a key challenge: the control features extracted for different modalities (e.g., Canny, Depth) reside in distinct and potentially incompatible feature spaces. As shown in Figure 4, we propose a unified version, SCALAR-Uni. The core idea is to map disparate control features into a common, modality-agnostic latent space. We utilize the image feature space itself as the target for this alignment, as it offers a rich and universal representation of visual concepts. We enforce this alignment by introducing an auxiliary loss term during training, which minimizes the L2 distance between the projected control representations and the corresponding image features:
@@
Figure 5: Visual results generated by SCALAR for class-to-image controllable generation.
Figure 6: Visualization of class-to-image controllable generation with SCALAR by changing class labels (e.g., “polecat” → “goldfish”).
Figure 7: Visualization of zero-shot class-to-image controllable generation task with SCALAR, including inpainting and outpainting. From top to bottom, Depth, Sketch, Normal, Canny, and HED are used as control image inputs.
@@ 4.1 Experimental Setup
Dataset. Our experiments are conducted on class-to-image (c2i) controllable generation. We conduct experiments on the ImageNet-256 [deng2009imagenet], using all 50K images from the validation set for evaluation. To assess the controllable generation capability of our model, we consider five types of conditions: Canny [canny1986computational], Depth [ranftl2020towards], Normal [vasiljevic2019diode], HED [xie2015holistically], and Sketch [su2021pixel].
Training Details. (1) We extract four features from the Image encoder $\mathcal{E}$ at depths $d-1-k\cdot\lfloor d/4\rfloor$ (k=0,1,2,3k{=}0,1,2,3). (2) We use DINOv2 [dinov2] as the image encoder for its strong representations and scalability. (3) For class-to-image (c2i) generation, we use pretrained VARs [var] with depths of 12, 16, 20, and 24. DINOv2-S is used for d12, and DINOv2-B otherwise, as detailed in Table 1. In this setting, our SCALAR and SCALAR-Uni are trained for 10 epochs with the AdamW optimizer on 8 NVIDIA H20 GPUs. (4) Following previous works [zhang2023adding, li2024controlnet++], the control block is zero-initialized. (5) For SCALAR-Uni, we sample different control conditions with equal probability.
Evaluation Metrics. We mainly employ two metrics: conditional consistency and image generation quality. We evaluate the conditional consistency by calculating the similarity between the input condition images and the extracted condition images from the generated images. Specifically, we use F1-Score to assess the similarity of Canny and Sketch, Root Mean Square Error (RMSE) for Normal and Depth, Structural Similarity Index Measure (SSIM) for HED and Lineart, and the mean Intersection-over-Union (mIoU) for Segmentation. For image generation quality, we use Fréchet Inception Distance (FID) [heusel2017gans] and Inception Score (IS) [salimans2016improved].
4.2 Class-to-image Experimental Results
We evaluate the class-to-image controllable generation performance of our proposed SCALAR and SCALAR-Uni on the ImageNet [deng2009imagenet]. As shown in Table 1, we compute both conditional consistency and image generation quality for the images generated by SCALAR and SCALAR-Uni. Furthermore, we compare our approach with existing controllable generation methods, including diffusion-based models [zhang2023adding, mou2024t2i], raster-scan AR models [li2024controlar], and VAR-based methods [controlvar, yao2024car].
@@ 4.2 Class-to-image Experimental Results
Figure 8: Visual results generated by our unified multi-condition control method SCALAR-Uni under varying control conditions for class-to-image controllable generation.
Results for SCALAR-Uni. Compared to other SoTA methods [zhang2023adding, mou2024t2i, li2024controlar, controlvar, yao2024car], SCALAR-Uni demonstrates clear advantages in image generation quality while maintaining high conditional consistency. When compared to SCALAR, SCALAR-Uni shows a slight drop in both generation quality and consistency, which may be attributed to the need to accommodate multiple control conditions, thereby reducing the effective training data per condition. Overall, SCALAR-Uni exhibits strong performance and generalization as a unified framework for multi-condition controllable generation. More visualization details can be found in Figure 8.
Figure 9: Visualization of zero-shot hybrid class-to-image controllable generation with SCALAR-Uni, combining two control types (e.g., Depth+Sketch) to generate images with both characteristics. Different class-condition injections are applied for each control pair.
@@ 4.3 Ablation Study
Ablations on Image Encoder $\mathcal{E}$. In Table 2, we conduct experiments using different image encoders (or pretraining schemes) towards different controls on ImageNet [deng2009imagenet], including Canny and Depth. Firstly, unlike previous approaches [li2024controlar, yao2024car], SCALAR and SCALAR-Uni adopt frozen image encoders [vit, dinov2, sam] to preserve the robust features obtained from large-scale self-supervised pretraining. In addition, inspired by prior work [ye2024biggait, bolya2025perception, ye2025biggergait], we extract multi-layer features including intermediate layers, instead of relying solely on the final output, which improves image generation quality while maintaining control consistency. As shown in Table 2 (c) to (g), DINOv2 [dinov2] outperforms other pretrained vision models such as ViT [vit] and SAM [sam]. Furthermore, comparing (d), (e), (c), and (g) indicates that increasing the scale of the image encoder significantly boosts performance under the same backbone. Considering the trade-off between parameter size and effectiveness, we use DINOv2-s for the d12 architecture and DINOv2-b for all other settings. Finally, the comparison from (g) to (j) demonstrates that deeper VAR networks lead to further improvements in both generation quality and control accuracy, suggesting that deeper models can better exploit visual representations.
@@
Ablations on Unified Control Alignment. We evaluate our SCALAR-Uni under different depths (d12, d16, d20, and d24) on ImageNet [deng2009imagenet]. While the model is trained with five types of control conditions, the effectiveness of Unified Control Alignment is validated only on Canny and Depth. As shown in Table 3, without the Unified Control Alignment, SCALAR-Uni is equivalent to SCALAR trained with mixed supervision across the five control types. Thanks to the strong general visual representation capability of DINOv2, SCALAR can effectively capture the differences across modalities and achieve preliminary alignment of multi-modal control signals. After incorporating Unified Control Alignment, SCALAR-Uni consistently outperforms SCALAR across different network depths in both image generation quality and control consistency. Results in Table 3 demonstrate that SCALAR-Uni, constructed by integrating SCALAR with Unified Control Alignment, offers a simple and effective solution for controllable image generation under diverse conditions.
@@ 5 Conclusion and Future Work
Future Work. (1) To align with the next-scale paradigm in VAR, we directly inject control features extracted from the image encoder into corresponding scales after resizing them via bilinear interpolation. While simple to implement, this approach is relatively coarse and lacks fine-grained semantic alignment, which may limit the further improvement of SCALAR’s controllability. This issue is worth further study. (2) SCALAR successfully combines the powerful generation capabilities of VAR with a simple and effective control injection mechanism. We are excited to inspire related tasks (e.g., image editing).
Table 2: Ablation of Image Encoder $\mathcal{E}$ and the backbone depth of VAR. $x$ cells mark Image Encoder parameters comparable to CAR [yao2024car] / ControlAR [li2024controlar], with a smaller backbone.